2^-3-3^-1 does that make the answer 1/2^3- 1/3^1. Thank you for your time
Answer by stanbon(75887)

Yes, that would be the 1st step.
Then 1/2^3 - 1/3^1
= 1/8 - 1/3
= (3-8)/24
= -5/24
Cheers,
stan H.